a golfer drives a ball toward a lake.The golf ball was struck at an angle of 27 above the horizontal with a speed of 5.6m/s.How far did the ball travel horizontally before hitting the water


1
Expert's answer
2021-10-24T18:25:40-0400

The range is given as follows (https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/):


"R = \dfrac{u^2\sin (2\theta)}{g}"

where "u = 5.6m\/s", "\theta = 27\degree" and "g = 9.8m\/s^2"

"R = \dfrac{5.6^2\cdot \sin(2\cdot 27\degree)}{9.8} \approx 2.6m"

Answer. 2.6 m.